In the following system of equations, what does x equal?
8x - 8y = 16
4x + 4y = -72

Answer by ewatrrr(24785)


Hi
solving System by Elimination
8x - 8y = 16
4x + 4y = -72
8x - 8y = 16
8x + 8y = -144
16x = -128
x = -8 and y = -10 |(-72-(-32))/4 = -40/4 |using 2nd Eq
CHECKING our Answer***
8x - 8y = 16
-64 + 80 = 16